determine how many bit strings of length can be formed, where three consecutive 0s are not allowed
1
Expert's answer
2021-09-07T13:50:06-0400

Let use the following example.

Determine how many bit strings of length 5 can be formed, where three consecutive 0s are not allowed.

The number of elements of the set of different binary strings of length 5 is "2^5=32" .

"\begin{matrix}

0& 0 & 0 & 0 & 0\\

0& 0 & 0 & 0 & 1\\

0& 0 &0 &1 &0 \\

0& 0 & 0 & 1 & 1

\end{matrix}"


"\begin{matrix}

1 & 0 & 0 & 0 & 0\\

1 & 0 & 0 & 0 & 1

\end{matrix}"


"\begin{matrix}

0 & 1 & 0 & 0 & 0\\

1& 1 & 0 & 0 & 0

\end{matrix}"


The number of bit strings of length 5 can be formed, where three consecutive 0s are not allowed, is 32-8=24